A man has 10 gallons of a 50% sulphuric acid mixture, and 20 gallons of a 20% solution, and 5 gallons of a 10% solution. He wants to use up all the 10% solution and make 15 gallons of 30% solution. How much of each solution should he use?
Answer by TimothyLamb(4379)

50x + 20y + 10z = 30(15)
50x + 20y + 10z = 450
---
z = 5
---
x + y + z = 15
---
system of linear equations:
50x + 20y + 10z = 450
x + y + z = 15
z = 5
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
get result from solver:
x= 6.66666667
y= 3.33333333
z= 5
---
answer:
x= 6 + 2/3 gal 50% solution
y= 3 + 1/3 gal 20% solution
z= 5 gal 10% solution
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php
---
Solve systems of linear equations to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php